# Ideal transition systems

### v2

@@ Abstract
This semi-expository article presents an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring. This method starts from a given set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish a particular transition recurrence among the chosen ideals. We describe explicit constructions of such systems in two motivating cases—namely, for the ideals of matrix Schubert varieties and their skew-symmetric analogues. Despite many formal similarities with these examples, for the symmetric versions of matrix Schubert varieties, it is an open problem to construct the same kind of transition system. We present several conjectures that would follow from such a construction, while also discussing the special obstructions arising in the symmetric case.
@@ 1 Introduction
Let $\mathsf{Mat}_{n\times n}$ be the variety of $n\times n$ matrices over an algebraically closed field $\mathbb{K}$. Write $\mathsf{B}_{n}\subset\mathsf{Mat}_{n\times n}$ for the group of invertible lower triangular matrices in this set.
A version of Gaussian elimination shows that every $M\in\mathsf{Mat}_{n\times n}$ can be written as $b_{1}wb_{2}^{\top}$ for $b_{1},b_{2}\in\mathsf{B}_{n}$ and a unique partial permutation matrix $w$. Equivalently, the $n\times n$ partial permutation matrices are representatives for the distinct $\mathsf{B}_{n}\times\mathsf{B}_{n}$-orbits on $\mathsf{Mat}_{n\times n}$. These orbits are called matrix Schubert cells $\mathsf{MSC}_{w}$, and their Zariski closures are called matrix Schubert varieties $\mathsf{MSV}_{w}$. We can express these using rank conditions as
$\mathsf{MSC}_{w}$ =$=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\},$ (1.1)
$\mathsf{MSV}_{w}$ =$=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\}.$
Here we write $[n]=\{1,2,\ldots,n\}$, and if $R,C\subseteq[n]$ then $M_{RC}$ denotes the submatrix of $M$ in rows $R$ and columns $C$. Thus, $M_{[i][j]}$ is the upper-left $i\times j$ corner of $M$.
@@
Fulton [13] showed that the equations setting these minors to zero also define $\mathsf{MSV}_{w}$ scheme-theoretically, i.e., that $I_{w}$ is a radical ideal (in fact, it is prime). Knutson and Miller [23] proved that the generating set of $I_{w}$ just described is actually a Gröbner basis under an appropriate term order. This fact allows for convenient computation of certain cohomological invariants associated to $I_{w}$, and can also be used to derive Fulton’s result.
In [29], we reproved these results using a new inductive approach. The key step is establishing a transition recurrence of the form
@@
There is a geometric reason for such recurrences to exist. If $w$ and $y$ are two partial permutation matrices, then $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is a closed $\mathsf{B}_{n}$-stable set and therefore decomposes as a union $\bigcup_{v\in\Phi}\mathsf{MSV}_{v}$. A result of Ramanathan [32] implies that the scheme $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is actually a reduced union of matrix Schubert varieties, so that the stronger equality $I_{w}+I_{y}=\bigcap_{v\in\Phi}I_{v}$ also holds. The transition recurrence then follows by an appropriate choice of $y$.
@@
In our previous work [29], we constructed Gröbner bases for the prime ideals of the skew-symmetric matrix Schubert varieties $\mathsf{MSV}^{\mathrm{ss}}_{w}$. These generating sets consist of Pfaffians of principal submatrices of a generic skew-symmetric matrix, and are more complicated to describe than the minors generating the ideals $I_{w}$ above. However, the proof technique based on transition recurrences is the same as in the classical case. It would be interesting to know if there is some geometric result similar to [32] that explains these skew-symmetric transition recurrences; at present, this is an open question.
The first, partially expository goal of this article is to present a streamlined version of the proof techniques just described. Specifically, we formalize an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring, in terms of what we call transition systems. Such a system is composed of a set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish transition recurrences analogous to (1.2) among the chosen ideals. Section 3 develops the general setup and nontrivial properties of transition systems, following some preliminaries in Section 2. Our motivating examples appear in Sections 4 and 5, which describe transition systems that compute Gröbner bases for the matrix Schubert varieties $\mathsf{MSV}_{w}$ and $\mathsf{MSV}^{\mathrm{ss}}_{w}$.
Our second goal in this paper is to explain some conjectures related to another interesting variant of the matrix Schubert construction, where $\mathsf{Mat}_{n\times n}$ is replaced by the subset of symmetric matrices $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. Using the same action as in the skew-symmetric case, one may consider the $\mathsf{B}_{n}$-orbits $\mathsf{MSC}^{\mathrm{sym}}_{w}$ in $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ and their closures $\mathsf{MSV}^{\mathrm{sym}}_{w}$, which we call symmetric matrix Schubert varieties. These may again be defined by the rank conditions in (1.1), but restricted to symmetric matrices. When $\mathbb{K}$ does not have characteristics two (as we assume in Example 1.2 below), the corresponding orbit representatives $w$ are provided by all symmetric $n\times n$ partial permutation matrices.
@@
Because of obstructions like this one, the set of ideals $I^{\mathrm{sym}}_{w}$ is too small to be assembled into a transition system as defined in Section 3. Nevertheless, we conjecture that a larger transition system exists which includes these ideals. We explain this conjecture precisely in Section 6 and present some evidence that it holds.
There, we also outline some appealing consequences of such a result, for which we have more computational support. For example, our main Conjecture 6.3 would imply that each ideal $I^{\mathrm{sym}}_{w}$ is both radical and prime, and that the ideal’s original generating set of minors is a Gröbner basis under an appropriate term order. These properties have been checked by computer when $\operatorname{\mathsf{char}}(\mathbb{K})=0$ for $n\leq 7$.
@@ Acknowledgments
This work was partially supported by Hong Kong RGC grants 16306120 and 16304122. We thank Darij Grinberg for helpful comments regarding Lemma 2.9.
@@ 2.1 Gröbner bases
The lexicographic term order on $\mathbb{K}[{\bf x}]$ has $x_{1}^{a_{1}}x_{2}^{a_{2}}\cdots x_{N}^{a_{N}}\leq x_{1}^{b_{1}}x_{2}^{b_{2}}\cdots x_{N}^{b_{N}}$ if and only if $(a_{1},a_{2},\ldots,a_{N})\leq(b_{1},b_{2},\ldots,b_{N})$ in lexicographic order.
@@
Later we will consider the reverse lexicographic term order on polynomial rings generated by variables $x_{ij}$ indexed by certain pairs $(i,j)\in\mathbb{Z}\times\mathbb{Z}$. In this context, we order the indexing pairs $(i,j)$ lexicographically to write products of variables in the form (2.1). Then the reverse lexicographic term order is the unique the total order $<$ such that:
@@
Fix a term order and suppose $f=\sum_{\mathsf{mon}}c_{\mathsf{mon}}\cdot\mathsf{mon}\in\mathbb{K}[{\bf x}]$ where the sum is over all monomials $\mathsf{mon}\in\mathbb{K}[{\bf x}]$, with each $c_{\mathsf{mon}}\in\mathbb{K}$. The initial term $\operatorname{init}(f)$ of $f$ is either the maximal monomial $\mathsf{mon}$ such that $c_{\mathsf{mon}}\neq 0$, or zero when $f=0$.
@@
Not every generating set for an ideal is a Gröbner basis. The ideal $I=\langle x_{1}x_{2}\rangle\subset\mathbb{K}[x_{1},x_{2}]$ is generated by $\{x_{2}-x_{1},x_{2}\}$, but this is not a Gröbner basis for any term order with $x_{1}<x_{2}$. However, if $I\subseteq\mathbb{K}[{\bf x}]$ is any ideal then $\operatorname{init}(\operatorname{init}(I))=\operatorname{init}(I)$. This means that if $S$ is any set of monomials in $\mathbb{K}[{\bf x}]$ then $S$ is a Gröbner basis for the ideal $\langle S\rangle$.
@@ 2.2 More on initial ideals
We continue the setup of the previous section. There are a few additional properties of initial ideals that will be of use later. To begin, we recall a few basic facts included in [29, Prop. 2.4]:
Proposition 2.7.
@@
The next result is a stronger version of [29, Lem. 2.5].
@@
The following proof was explained to us by Darij Grinberg. Define a bridge to be a pair $(f,g)$ where $f\in I$ and $g\in J$ and $\operatorname{init}(f)=\operatorname{init}(g)>\operatorname{init}(f-g)$, where $>$ denotes our term order on monomials and where $\mathsf{mon}>0$ for all monomials. Note that this requires $f$ and $g$ to have the same leading monomial with the same coefficient, unless one has $f=g=0$. Define the floor of a bridge $(f,g)$ to be $\operatorname{init}(f-g)$.
@@
First suppose $\operatorname{init}(f-g)\in\operatorname{init}(I)$. Then $\operatorname{init}(f-g)=\operatorname{init}(i)$ for some $0\neq i\in I$. Define $\lambda\in\mathbb{K}$ to be the leading term of $f-g$ divided by the leading term of $i$. Then the pair $(f-\lambda i,g)$ is a bridge with a lower floor than $(f,g)$. By induction, we have $\operatorname{init}(f-\lambda i)=\operatorname{init}(g)\in\operatorname{init}(I\cap J)$. Since $\operatorname{init}(f)=\operatorname{init}(f-\lambda i)$, this proves the claim when $\operatorname{init}(f-g)\in\operatorname{init}(I)$.
The argument when $\operatorname{init}(f-g)\in\operatorname{init}(J)$ is similar. For this case, one checks that if $0\neq j\in J$ is an element with $\operatorname{init}(f-g)=\operatorname{init}(j)$ and $\mu=\frac{\operatorname{init}(f-g)}{\operatorname{init}(j)}\in\mathbb{K}$, then the pair $(f,g+\mu j)$ is a bridge with lower floor than $(f,g)$, so by induction $\operatorname{init}(f)=\operatorname{init}(g+\mu j)=\operatorname{init}(g)\in\operatorname{init}(I\cap J)$.
@@
The following technical property repeats a result from [29].
Lemma 2.10 ([29, Lem. 2.6]).
@@ 3.1 Main definition
Throughout, we let $A$ be an affine variety over an algebraically closed field $\mathbb{K}$. Let $\mathsf{Index}$ be a finite indexing set, choose variables $x_{i}$ to identify $\mathbb{K}[A]=\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ with a ring of polynomials, and then fix a term order on this ring.
For an ideal $I\subseteq\mathbb{K}[A]$ and $0\neq f\in\mathbb{K}[A]$ let $(I:\langle f\rangle)=\tfrac{1}{f}(I\cap\langle f\rangle)=\{g\in\mathbb{K}[A]:fg\in I\}.$ This ideal quotient is an ideal satisfying $I\subseteq(I:\langle f\rangle)$, with equality if and only if $I\cap\langle f\rangle=fI$. Notice that $(I:\langle f\rangle)=\mathbb{K}[A]$ if and only if $f\in I$. Hence, if $I$ is maximal with $f\notin I$, then $I=(I:\langle f\rangle)$.
@@
    If $I^{\prime}:=(I:\langle f\rangle)$ is not equal to $I$ and $J^{\prime}:=(J:\langle\operatorname{init}(f))\rangle)$ then $(I^{\prime},J^{\prime})\in\mathcal{T}$.
@@
We claim that $\mathcal{T}=\{(I_{S},J_{S}):S\in\mathscr{S}\}$ is a transition system for any subset-closed family $\mathscr{S}$ of nonempty finite subsets of Z$\mathbb{Z}_{>0}$. The ideal $I_{S}$ is maximal when $|S|=1$, and then $\operatorname{init}(I_{S})=\langle x\rangle=J_{S}.$ If $|S|>1$ and $m\in S$ is any element, then conditions (T3) and (T4) in Definition 3.1 hold for
@@
If $I$ is maximal then $J=\operatorname{init}(I)$ by assumption (T1), while if $I=J$ then $I$ is a monomial ideal so $J=I=\operatorname{init}(I)$. Suppose $I$ is not maximal and $I\neq J$. Assume $f\notin I$ and $\varnothing\subsetneq\Phi\subseteq\mathcal{T}$ satisfy conditions (T3) and (T4) in Definition 3.1. Then $I\neq I+\langle f\rangle\subseteq\bigcap_{(P,Q)\in\Phi}P,$ so
@@
Since $\mathbb{K}[A]$ is Noetherian and every $(P,Q)\in\Phi$ has $I\subsetneq I+\langle f\rangle\subseteq P$, we may assume by induction that $\operatorname{init}(P)=Q$ for all $(P,Q)\in\Phi$. As we have $J\subseteq\operatorname{init}(I)$, it follows that
@@
Suppose $(I,J)$ belongs to a transition system $\mathcal{T}$. Assume we have a non-constant polynomial $f\notin I$ and a finite set $\varnothing\subsetneq\Phi\subseteq\mathcal{T}$ satisfying (T3) and (T4) in Definition 3.1. Then:
@@
For a given set $\mathscr{I}$ of ideals in $\mathbb{K}[A]$, there exists at most one transition system $\mathcal{T}$ with =$\mathscr{I}=\{I:(I,J)\in\mathcal{T}\}$.
@@
If $(I,J)\in\mathcal{T}$ then $\operatorname{init}(I)=J$ by Theorem 3.3, so $G_{I}$ is a Gröbner basis by Corollary 2.8. ∎

### v3

@@ Abstract
We study an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring. This method starts from a given set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish a particular transition recurrence among the chosen ideals. We describe explicit constructions of such systems in two motivating cases—namely, for the ideals of matrix Schubert varieties and their skew-symmetric analogues. Despite many formal similarities with these examples, for the symmetric versions of matrix Schubert varieties, it is an open problem to construct the same kind of transition system. We present several conjectures that would follow from such a construction, while also discussing the special obstructions arising in the symmetric case.
@@ 1 Introduction
Fix a positive integer $n$ and let $\mathsf{Mat}_{n\times n}$ be the variety of $n\times n$ matrices over an algebraically closed field $\mathbb{K}$. Write $\mathsf{B}_{n}\subset\mathsf{Mat}_{n\times n}$ for the group of invertible lower triangular matrices in this set.
Recall that a partial permutation matrix is a binary matrix with at most one nonzero entry in each row and column. A version of Gaussian elimination shows that every $M\in\mathsf{Mat}_{n\times n}$ can be written as $b_{1}wb_{2}^{\top}$ for $b_{1},b_{2}\in\mathsf{B}_{n}$ and a unique partial permutation matrix $w$. Equivalently, the $n\times n$ partial permutation matrices are representatives for the distinct $\mathsf{B}_{n}\times\mathsf{B}_{n}$-orbits on $\mathsf{Mat}_{n\times n}$. These orbits are called matrix Schubert cells $\mathsf{MSC}_{w}$, and their Zariski closures are called matrix Schubert varieties $\mathsf{MSV}_{w}$. We can express these sets using rank conditions as
$\mathsf{MSC}_{w}$ $=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\},$ (1.1)
$\mathsf{MSV}_{w}$ $=\{M\in\mathsf{Mat}_{w}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $i,j\in[n]$}\}.$
Here we write $n]=\{1,2,\ldots,n\}$, and if $R,C\subseteq[n]$ then $M_{RC}$ denotes the submatrix of $M$ in rows $R$ and columns $C$. Thus, $M_{[i][j]}$ is the upper-left $i\times j$ corner of $M$.
@@
Fulton [13] showed that the equations setting these minors to zero also define $\mathsf{MSV}_{w}$ scheme-theoretically, i.e., that $I_{w}$ is a radical ideal (in fact, it is prime). Knutson and Miller [25] proved that this generating set of minors is actually a Gröbner basis for $I_{w}$ under an appropriate term order. This fact allows for convenient computation of certain cohomological invariants associated to $I_{w}$, and can also be used to derive Fulton’s result.
In [31], we reproved these results using a new inductive approach. The key step in this approach was establishing a transition recurrence of the form
@@
There is a geometric reason for such recurrences to exist. If $w$ and $y$ are two partial permutation matrices, then $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is a closed $\mathsf{B}_{n}$-stable set and therefore decomposes as a union $\bigcup_{v\in\Phi}\mathsf{MSV}_{v}$. A result of Ramanathan [34] implies that the scheme $\mathsf{MSV}_{w}\cap\mathsf{MSV}_{y}$ is actually a reduced union of matrix Schubert varieties, so the stronger equality $I_{w}+I_{y}=\bigcap_{v\in\Phi}I_{v}$ also holds. The transition recurrence then follows by an appropriate choice of $y$.
@@
In our previous work [31], we constructed Gröbner bases for the prime ideals of the skew-symmetric matrix Schubert varieties $\mathsf{MSV}^{\mathrm{ss}}_{w}$. These generating sets consist of Pfaffians of principal submatrices of a generic skew-symmetric matrix, and are more complicated than the minors generating the ideals $I_{w}$ above. However, the proof technique based on transition recurrences is the same as in the classical case. It would be interesting to know if there is some geometric result similar to [34] that explains these skew-symmetric transition recurrences; at present, this is an open question.
The first, partially expository goal of this article is to present a streamlined version of the proof techniques just described. Specifically, we formalize an inductive method of computing initial ideals and Gröbner bases for families of ideals in a polynomial ring, in terms of what we call transition systems. Such a system is composed of a set of pairs $(I,J)$ where $I$ is any ideal and $J$ is a monomial ideal contained in the initial ideal of $I$. These containments become a system of equalities if one can establish transition recurrences analogous to (1.2) among the chosen ideals. Section 3 develops the general setup and nontrivial properties of transition systems, following some preliminaries in Section 2. Our motivating examples appear in Sections 4 and 5, where we present transition systems that compute Gröbner bases for the matrix Schubert varieties $\mathsf{MSV}_{w}$ and $\mathsf{MSV}^{\mathrm{ss}}_{w}$.
Our second goal in this paper is to explain some conjectures related to another interesting variant of the matrix Schubert construction, where $\mathsf{Mat}_{n\times n}$ is replaced by the subset of symmetric matrices $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. Using the same action as in the skew-symmetric case, one may consider the $\mathsf{B}_{n}$-orbits $\mathsf{MSC}^{\mathrm{sym}}_{w}$ in $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$ and their closures $\mathsf{MSV}^{\mathrm{sym}}_{w}$, which we call symmetric matrix Schubert varieties. These may again be defined by the rank conditions in (1.1), but restricted to symmetric matrices. When $\mathbb{K}$ does not have characteristic two (as we assume in Example 1.2 below), the corresponding orbit representatives $w$ are provided by all symmetric $n\times n$ partial permutation matrices.
@@
Because of obstructions like this one, the set of ideals $I^{\mathrm{sym}}_{w}$ is too small to be assembled into a transition system as defined in Section 3. Nevertheless, we conjecture that a larger transition system exists which includes these ideals. We explain this conjecture precisely in Section 6 (see Conjecture 6.3) and present some evidence that it holds.
There, we also outline some appealing consequences of such a result, for which we have more computational support. In detail, our main conjecture would imply that each ideal $I^{\mathrm{sym}}_{w}$ is both radical and prime (Conjecture 6.7), and that the ideal’s original generating set of minors is a Gröbner basis under an appropriate term order (Conjecture 6.6). These properties have been checked by computer when $\operatorname{\mathsf{char}}(\mathbb{K})=0$ for $n\leq 7$. Conjecture 6.3 would also imply that $I^{\mathrm{sym}}_{w}$ is the ideal of the variety $\mathsf{MSV}^{\mathrm{sym}}_{w}$ (Conjecture 6.5).
In Section 7 we study a more involved application of this last conjectural property. Specifically, assuming Conjecture 6.5 as a hypothesis, we show that the orthogonal Grothendieck polynomials introduced in [28] have well-defined stable limits. This reduces the open problem [28, Prob. 5.3] to the task of constructing a transition system that includes all of the ideals $I^{\mathrm{sym}}_{w}$.
@@ Acknowledgments
This work was partially supported by Hong Kong RGC grants 16306120 and 16304122. We thank Darij Grinberg for helpful comments regarding Lemma 2.9, and Allen Knutson for helping to clarify our approach in §7.
@@ 2.1 Gröbner bases
The lexicographic term order on $\mathbb{K}[{\bf x}]$ has $x_{1}^{a_{1}}x_{2}^{a_{2}}\cdots x_{N}^{a_{N}}\leq x_{1}^{b_{1}}x_{2}^{b_{2}}\cdots x_{N}^{b_{N}}$ if and only if $(a_{1},a_{2},\ldots,a_{N})\leq(b_{1},b_{2},\ldots,b_{N})$ in lexicographic order, meaning that if there is a smallest index $i$ with $a_{i}\neq b_{i}$ then $a_{i}<b_{i}$.
@@
[…]
@@
Fix a term order and suppose $f=\sum_{\mathsf{mon}}c_{\mathsf{mon}}\cdot\mathsf{mon}\in\mathbb{K}[{\bf x}]$ where the sum is over all monomials $\mathsf{mon}\in\mathbb{K}[{\bf x}]$ and all but finitely many of the coefficients $c_{\mathsf{mon}}\in\mathbb{K}$ are zero. The initial term $\operatorname{init}(f)$ of $f$ is either the maximal monomial $\mathsf{mon}$ such that $c_{\mathsf{mon}}\neq 0$, or zero when $f=0$.
@@
Not every generating set for an ideal is a Gröbner basis. The ideal $I=\langle x_{1},x_{2}\rangle\subset\mathbb{K}[x_{1},x_{2}]$ is generated by $\{x_{2}-x_{1},x_{2}\}$, but this is not a Gröbner basis for any term order with $x_{1}<x_{2}$. However, if $I\subseteq\mathbb{K}[{\bf x}]$ is any ideal then $\operatorname{init}(\operatorname{init}(I))=\operatorname{init}(I)$. This means that if $S$ is any set of monomials in $\mathbb{K}[{\bf x}]$ then $S$ is a Gröbner basis for the ideal $\langle S\rangle$.
@@ 2.2 More on initial ideals
We continue the setup of the previous section. There are a few additional properties of initial ideals that will be of use later. To begin, we recall a few basic facts derived in [31].
Proposition 2.7 (See [31, Prop. 2.4]).
@@
The next result is a stronger version of [31, Lem. 2.5].
@@
The following proof was explained to us by Darij Grinberg. Define a bridge to be a pair $(f,g)$ where $f\in I$ and $g\in J$ and $\operatorname{init}(f)=\operatorname{init}(g)>\operatorname{init}(f-g)$, where >> denotes our term order on monomials and where $\mathsf{mon}>0$ for all monomials. Note that this requires $f$ and $g$ to have the same leading monomial with the same coefficient, unless one has $f=g=0$. Define the floor of a bridge $(f,g)$ to be $\operatorname{init}(f-g)$.
@@
First suppose $\operatorname{init}(f-g)\in\operatorname{init}(I)$. Then $\operatorname{init}(f-g)=\operatorname{init}(i)$ for some $0\neq i\in I$. Define $\lambda\in\mathbb{K}$ to be the leading term of fgf-g divided by the leading term of $i$. Then the pair $(f-\lambda i,g)$ is a bridge with a lower floor than $(f,g)$. By induction, we have $\operatorname{init}(f-\lambda i)=\operatorname{init}(g)\in\operatorname{init}(I\cap J)$. Since $\operatorname{init}(f)=\operatorname{init}(f-\lambda i)$, this proves the claim when $\operatorname{init}(f-g)\in\operatorname{init}(I)$. The argument when $\operatorname{init}(f-g)\in\operatorname{init}(J)$ is similar.
@@
Lemma 2.10 ([31, Lem. 2.6]).
@@ 3.1 Main definition
Throughout, we let $A$ be an affine space over an algebraically closed field $\mathbb{K}$. Let $\mathsf{Index}$ be a finite indexing set, choose variables $x_{i}$ to identify $\mathbb{K}[A]=\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ with a ring of polynomials, and then fix a term order on this ring.
For an ideal $I\subseteq\mathbb{K}[A]$ and a nonzero element $0\neq f\in\mathbb{K}[A]$ let
$(I:\langle f\rangle)=\tfrac{1}{f}(I\cap\langle f\rangle)=\{g\in\mathbb{K}[A]:fg\in I\}.$ (3.1)
This ideal quotient is an ideal satisfying $I\subseteq(I:\langle f\rangle)$, with equality if and only if $I\cap\langle f\rangle=fI$. Since $(I:\langle f\rangle)=\mathbb{K}[A]$ if and only if $f\in I$, it follows that if $I$ is maximal with $f\notin I$ then $I=(I:\langle f\rangle)$.
@@
    If $I^{\prime}:=(I:\langle f\rangle)$ is not equal to $I$ and $J^{\prime}:=(J:\langle\operatorname{init}(f)\rangle)$ then $(I^{\prime},J^{\prime})\in\mathcal{T}$.
@@
[…]
@@
If $I$ is maximal then $J=\operatorname{init}(I)$ by assumption (T1), while if $I=J$ then $I$ is a monomial ideal so $J=I=\operatorname{init}(I)$. Suppose $I$ is not maximal and $I\neq J$. Assume $f\notin I$ and $\varnothing\subsetneq\Phi\subseteq\mathcal{T}$ satisfy conditions (T3) and (T4) in Definition 3.1. Then $I\neq I+\langle f\rangle\subseteq\bigcap_{(P,Q)\in\Phi}P$ so
@@
Since $\mathbb{K}[A]$ is Noetherian and every $(P,Q)\in\Phi$ has $I\subsetneq I+\langle f\rangle\subseteq P$, we may assume by induction that $\operatorname{init}(P)=Q$ for all $(P,Q)\in\Phi$. As we have $J\subseteq\operatorname{init}(I)$ it follows that
@@
Suppose $(I,J)$ belongs to a transition system $\mathcal{T}$. Assume $f\notin I$ is a non-constant polynomial and $\Phi\subseteq\mathcal{T}$ is a nonempty finite set satisfying (T3) and (T4) in Definition 3.1. Then:
@@
For a given set $\mathscr{I}$ of ideals in $\mathbb{K}[A]$, there exists at most one transition system $\mathcal{T}$ with $\mathscr{I}=\{I:(I,J)\in\mathcal{T}\}$.
@@
If $(I,J)\in\mathcal{T}$ then $\operatorname{init}(I)=J$ by Theorem 3.3 so $G_{I}$ is a Gröbner basis by Corollary 2.8. ∎